can you please help me with this problem
same direction:
a truck left the terminal traveling 50mi/h.
a heavier truck traveled at 60mph. it left 2 hours later.
they both follow the same route.
how many hours will it take truck #1 to catch up with truck #2

Answer by ewatrrr(24785)


Hi,
D = r*t This question sets up that the distance traveled is the same for each.
Question states: truck traveling at 60mph...left 2 hours later
Let (t + 2)and t represent the respective times traveled by the trucks
50mph(t+2) = 60mph*t
solving for t
50t + 100 = 60t
100 = 10t
10hr = t, time it will take for the truck traveling 60mph to catch up
CHECKING OUR ANSER
50mph*12hr = 60mph*10hr = 600mi (misc info as to the distance traveled by each)